OpenStudy (anonymous):

describe x^3 as a graph

OpenStudy (anonymous):

\(As\ x \rightarrow -\infty, y \rightarrow -\infty\\As\ x \rightarrow +\infty, y\rightarrow +\infty\) and it is symmetry at the origin.

OpenStudy (anonymous):

how would you call the swerved image?